OpenStudy (wesdg1978):

Using f(x)=(x+3)/X on [1,6] I'm supposed to find the point c guaranteed by the Mean Value Theorem. For f(6)=(6+3)/6=3/2 and f(1)=(1+3)/1=4 This gives f '(c)=(3/2-4)/(6-1) = (-5/2)/5=-1/2 What am I supposed to do now? I took the derivative of f(x) and got -3/x^2=0 Am I supposed to do something with that or what? Any hints/help is greatly appreciated!

OpenStudy (amistre64):

so there is at least one c in [a,b] such that f'(c) is equal to the slope between f(a) and f(b)

OpenStudy (wesdg1978):

yes,

OpenStudy (amistre64):

(x+3) (x)^-1 (x)^-1 - (x+3) (x)^-2 3/x^2 = -1/2; check to make sure your slope is correct between the end points

OpenStudy (amistre64):

3/c^2 = k 3/k = c^2 sqrt(3/k) = c , between 1 and 6

OpenStudy (wesdg1978):

I got -3/x^2

OpenStudy (amistre64):

:) fine -3/x^2 then

OpenStudy (wesdg1978):

lol

OpenStudy (amistre64):

sqrt(6) seems to be good for x tho

OpenStudy (amistre64):

or c

OpenStudy (wesdg1978):

OK, so you don't set it equal to 0 you set it equal to the f '(c) which was -1/2 and then that gives the point on the interval

OpenStudy (amistre64):

f' defines the slope at a single point; the mean value thrm states that there is some intermediary point between the ends that equals the slope between the ends if that slope is not 0 then I would not equate it to zero

OpenStudy (wesdg1978):

Yeah, I don't know why my original thought was to set it equal to zero, it just seems to be "the thing to do" in math. :)

OpenStudy (amistre64):

min/max is set to 0 ;) you did fine tho

OpenStudy (wesdg1978):

Just a technicality question, would the answer technically be plus or minus the sqrt(6) but it can't be minus because the interval is on [1,6] ?

OpenStudy (amistre64):

since -sqrt6 is not in the interval, you cant really say that its a solution within the stated interval

OpenStudy (wesdg1978):

Right, that's what I was getting at. If the interval was say [-6,6] then you would have to include minus sqrt(6) though right?

OpenStudy (amistre64):

not really, the interval needs to be continuous; and this is not continuous at x=0

OpenStudy (amistre64):

granted it works out fine in this case, but in general thats a bad idea

OpenStudy (amistre64):

the slope of the line between -6 and 6 is prolly zero

OpenStudy (wesdg1978):

OK, thanks for your help!

OpenStudy (amistre64):

 its very close to zero along -6 to 6 ... and since its discontinuous along the interval you cant guarantee a solution